OpenStudy (cicilybailey14):

So confused...will Medal!!! Find the range of the function s(t) = t^2 - 1 given the domain { -1, 0.5, 3.7 } { -0.75, -2, 12.69 } { -2, 0, 12.69 } { -0.75, 0, 12.69 } { -0.75, -2, 6.4 }

OpenStudy (danjs):

The range will be the numbers you get when you put in t = -1 t = 0.5 and t = 3.7 into s(t) = t^2 - 1

OpenStudy (danjs):

for example t = -1 s(-1) = (-1)^2 - 1 s(-1) = 1 - 1 s(-1) = 0

OpenStudy (danjs):

So zero is one of the three numbers, do the same for the other two t values

OpenStudy (cicilybailey14):

@DanJS C?

OpenStudy (danjs):

um..

OpenStudy (danjs):

yep!

OpenStudy (cicilybailey14):

it's either B or C lol

OpenStudy (cicilybailey14):

oh okay haha

OpenStudy (danjs):

i didnt calculate it, i just had to

OpenStudy (cicilybailey14):

it fine! I would fail without your help lmao!

OpenStudy (danjs):

for that kind of problem, just plug in the domain values into the function, it will give you the range values